17. Solve the system using the addition method. -5x - 9y = 28 - 4x + 5y = -2 O A.(-2,-2) O B. (7,-4) O C.(7,6) O D.(-11,3)
Answer #1

solve the system using the Addition methool -5X-94 = 28 - -4x+sy = -2 - -4x(i) → 20x +367 =-1/2 - () sxli) -20% +25y = -10 -